# OpenFGL: A Comprehensive Benchmarks for Federated Graph Learning

### v1

@@ 4.4. Efficiency Evaluation
To answer Q8, we provide the efficiency reports in Fig. 2 and obtain the following observations: (1) Prototype-based methods such as FedProto, FedTGP, and FGGP exhibit significant advantages in terms of less communication cost by transmitting prototype representations, yet they require additional computation either on the client or server side to achieve satisfactory performance, thus losing the advantage in terms of running time. (2) Cross-client collaborative methods such as FedGL and FedSage+ encounter reduced deployment efficiency due to added delays resulting from inter-client communication and synchronization. (3) Decoupled approaches such as AdaFGL, which aim to maximize local computational capacity while minimizing communication costs, demonstrate significant efficiency advantages. Consequently, we conclude that C10: FGL algorithms leveraging prototypes and decoupled techniques (i.e., multi-client collaboration then local updates) demonstrate substantial potential in applications with stringent efficiency requirements.
4.5. OpenFGL Algorithm Tutorial
In this section, we elucidate the design principles of the OpenFGL library from the perspective of algorithm design, which offers a unified API encompassing all FL/FGL algorithms. This approach enhances future technological advancement in a user-centric manner. Overall, our proposed OpenFGL employs the FGLTrainer Class to manage communication in each training round between clients and servers. This entails the exchange of messages at either the clients or server, thereby promoting federated collaboration. Specifically, the FGLTrainer Class utilizes a publicly available dictionary-type variable, message_pool, to aggregate messages from both clients and servers. In other words, OpenFGL directs all communications into this publicly available message_pool. Subsequently, users are encouraged to customize the FGLClient Class and FGLServer Class to modify the content and reception rules of messages, tailored to the specific algorithms. To provide an intuitive and reader-friendly description, we present implementations of PyTorch-style algorithms for FGLTrainer. We exemplify this by using FedAvgClient and FedAvgServer as instances of the FedAvg algorithm as follows:
# FGL Trainer class for client-server collaboration
@@ 4.5. OpenFGL Algorithm Tutorial
fgl_dataset = FGLDataset(args)
self.device = torch.device(f"cuda:{args.gpuid}" if (torch.cuda.is_available() and args.use_cuda) else "cpu")
self.clients = [load_client(args, client_id, fgl_dataset.local_data[client_id], fgl_dataset.processed_dir, self.message_pool, self.device) for client_id in range(self.args.num_clients)]
self.server = load_server(args, fgl_dataset.global_data, fgl_dataset.processed_dir, self.message_pool, self.device)
# Evaluation and history logger
self.evaluation_result = {"best_round":0}
if self.args.task in ["graph_cls", "graph_reg", "node_cls", "link_pred"]:
for metric in self.args.metrics:
[…]
elif self.args.task in ["node_clust"]:
for metric in self.args.metrics:
[…]
self.logger = Logger(args, self.message_pool, fgl_dataset.processed_dir)
@@
# Randomly activate partial clients
sampled_clients = sorted(random.sample(list(range(self.args.num_clients)), int(self.args.num_clients * self.args.client_frac)))
[…]
# Update message pool
@@
FGLTrainer. This class manages the flows of message and command among multiple clients and a central server in FGL processes. It begins by initializing the device, data, clients, and server using configuration parameters. Subsequently, in each specific federated training method, the trainer coordinates multiple communication rounds. During each round, it first selects a designated set of clients, then updates a message pool to reflect communication status, and finally dispatches server messages to the selected clients. Upon receiving the message, clients execute tasks locally and transmit updates to the server for global processing. Notably, Each communication round contains a evaluation phase. The Pytorch-style implementation of FGLTrainer is depicted in Algorithm 1.
# Client for each FL/FGL algorithm must inherit from the base class BaseClient
@@
def __init__(self, args, client_id, data, data_dir, message_pool, device):
super(FedAvgClient, self).__init__(args, client_id, data, data_dir, message_pool, device)
@@
for (local_param, global_param) in zip(self.task.model.parameters(), self.message_pool["server"]["weight"]):
@@
# Local training supports a variety of downstream tasks
@@
"num_samples": self.task.num_samples,
"weight": list(self.task.model.parameters())
@@
FedAvgClient. The implementation of this class within the OpenFGL framework is notably straightforward. Users only need to customize two modules: (1) Implement the local execution, where FedAvg involves downloading the global model and performing local training; (2) Define the messages that the client sends to the server, where FedAvg includes the local sample count and the local model weights. The Pytorch-style implementation of FedAvgClient is provided in Algorithm 2.
# Client for each FL/FGL algorithm must inherit from the base class BaseClient
@@
def __init__(self, args, global_data, data_dir, message_pool, device):
super(FedAvgServer, self).__init__(args, global_data, data_dir, message_pool, device)
@@
# Receive messages (number of samples, local model weight) from each client stored in the message pool to conduct global model aggregation
@@
[…]
for it, client_id in enumerate(self.message_pool["sampled_clients"]):
[…]
if it == 0:
global_param.data.copy_(weight * local_param)
else:
global_param.data += weight * local_param
@@
"weight": list(self.task.model.parameters())
Algorithm 3 OpenFGL-FGLServer: FedAvg Example.Pytorch style.
FedAvgServer. Similarly to FedAvgClient, users need to customize two algorithm-specific modules: (1) Implement the global execution process (i.e., FedAvg: global weighted model averaging based on the sample size); (2) Define the messages that the server sends to clients (i.e., FedAvg: the global model weights). The Pytorch-style implementation of FedAvgServer is illustrated in Algorithm 3.
@@ 5. Conclusion and Future Directions
In this paper, we first present a comprehensive overview of the current research progress in the FGL field and the significant potential of this technology for deployment in real-world graph-based database applications. Subsequently, we propose OpenFGL, a unified and comprehensive FGL benchmark, which encompasses 18 recently proposed SOTA FGL algorithms (Year: 2021-2024) and 38 datasets from 16 application domains for 5 graph-based downstream tasks across 2 prevalent FGL scenarios in the real world. The goal of our work is to fairly examine the current state of FGL development and offer key insights for future research endeavors. Leveraging OpenFGL, we conduct extensive experiments aimed at unveiling the performance of FGL algorithms from 3 orthogonal perspectives: effectiveness, robustness, and efficiency. Our investigations reveal promising advancements achieved by FGL studies but also highlight limitations in real-world deployments, such as vulnerability in inadequate node descriptions, robustness, and scalability. To inspire future research, we combine experimental Conclusions to present the following significant challenges and promising directions.
@@
For Robustness, (3) Personalized FGL (C5 and C7). In real-world scenarios, the robustness of FGL against client-specific noise and low client participation in communication is essential. During federated training, these factors significantly impact the attainment of global consensus, thereby hindering high-quality supervision provided for local training. Fortunately, personalized techniques can leverage local reliable knowledge to establish unbiased global consensus and customize broadcasts for each client, guiding local updates. This mitigates optimization challenges arising from noise data and limited data sources. Consequently, personalized FGL algorithms emerge as an effective strategy for addressing these concerns. (4) Multi-client Collaboration FGL (C6 and C8). The challenges related to sparsity and generalization often arise from expensive manual annotation and various application scenarios. During our investigation, we found that promoting server-side multi-client collaboration can extract global insights from sparse data for local training (e.g., federated knowledge-driven subgraph generation and augmentation in FedSage+ and AdaFGL). Additionally, this collaborative approach can capture shared semantic knowledge across data domains to facilitate robust generalization. Therefore, there is an emergency demand for more efficient multi-client collaboration FGL algorithms. (5) Privacy-preserve FGL (C9). The goal of FGL is to achieve multi-client collaborative training in a privacy-preserving manner without direct data sharing. However, current FGL algorithms, in pursuit of superior performance, increasingly share local information, raising potential concerns. Meanwhile, they are sensitive to additional DP-based privacy preservation. Therefore, the development of FGL algorithms with strict privacy requirements is imperative. Meanwhile, more privacy-preserving technologies (e.g., homomorphic encryption and secure multi-party computation) should be considered for further explorations.
@@
FGL should establish federated collaboration standards for various graph types (e.g., directed, signed, hypergraphs, heterogeneous) and learning paradigms (e.g., unsupervised, few-shot, continual, unlearning) based on the different downstream tasks. However, FGL remains a burgeoning field with numerous research gaps. Nevertheless, we are committed to continually enhancing OpenFGL to support future research endeavors. For instance, we are progressively refining the execution standards for federated heterogeneous graph learning. Notably, considering the space constraints and the need for a clear and reader-friendly presentation, we provide an overview and corresponding experimental results in  (Ope, 2024) (A.9).
@@ A.1. Dataset Description
In Graph-FL, we conduct experiments on the compounds networks (MUTAG, BZR, COX2, DHFR, PTC-MR, AIDS, NCI1, hERG) (Debnath et al., 1991; Sutherland et al., 2003; Helma et al., 2001; Riesen and Bunke, 2008; Wale et al., 2008; Gaulton et al., 2017), protein networks (ENZYMES, DD, PROTEINS) (Borgwardt et al., 2005; Dobson and Doig, 2003; Helma et al., 2001), collaboration network (COLLAB) (Leskovec et al., 2005), movie networks (IMDB-BINARY, IMDB-MULTI) (Yanardag and Vishwanathan, 2015), super-pixel networks (MNISTSuperPixels) (Monti et al., 2017), and point cloud networks (ShapeNet) (Yi et al., 2016). Notably, the node features in Graph-FL often consist of node attributes and labels. In most graph-level GNNs, these components are typically concatenated and fed into the model. Therefore, we report the feature dimension by concatenating node attributes and labels. Notably, super-pixel networks represent images as graphs, with nodes as super-pixels and edges capturing spatial or semantic relationships, enabling effective graph-based classification by leveraging image structure. Detailed dataset description is as follows:
@@
BZR (Sutherland et al., 2003) is a bioinformatics dataset used for compound activity prediction, with a primary focus on a collection of benzimidazole compounds. The dataset is designed to indicate the concentration of each compound necessary to inhibit the activity of specific biomolecules, providing valuable insights into the effectiveness of these compounds in biological processes.
COX2 (Sutherland et al., 2003) is a dataset centered on Cyclooxygenase-2, an enzyme that plays a critical role in inflammation and pain mechanisms. The dataset is utilized to classify various compounds and predict their potential inhibition potency against the COX-2 enzyme, which is a key target in drug development for anti-inflammatory therapies.
@@
ENZYMES (Borgwardt et al., 2005) is a comprehensive dataset containing 600 protein tertiary structures, meticulously curated from the BRENDA enzyme database. Within the ENZYMES dataset, researchers can explore the intricate structures of six unique enzymes, providing a rich resource for computational analysis and machine learning applications.
@@ A.2. Baseline Description
Regarding (2), while some prevalent FL approaches (FedAvg (McMahan et al., 2017), FedProx (Li et al., 2020), Scaffold (Karimireddy et al., 2020), MOON (Li et al., 2021), and FedDC (Gao et al., 2022)) have demonstrated effectiveness in computer vision-based (CV-based) federated scenarios, recent FGL studies contend that they only achieve sub-optimal performance. This weakness is attributed to their failure to incorporate topological information during the collaborative optimization process of FGL. Therefore, we only implement prevalent FL as baselines to assist future FGL developers in evaluating the effectiveness of their proposed methods.
@@
Building upon this foundation, to provide effective baselines, we have particularly implemented prototype-based methods (FedProto (Tan et al., 2022), FedNH (Dai et al., 2023), FedTGP (Zhang et al., 2024a)). This has sparked a research trend in recent CV-based FL, as this technology can comprehensively address the aforementioned multi-level heterogeneity challenges. The prevalent graph-independent FL baseline details implemented in our proposed OpenFGL are listed below:
FedAvg (McMahan et al., 2017) serves as a foundational method in FL, enabling decentralized model training across diverse devices while preserving data privacy. Initiated by a central server that distributes a global model, clients independently execute local updates through stochastic gradient descent. Subsequently, these updates are aggregated by the server via averaging to refine the global model, with the cycle repeating until convergence.
FedProx (Li et al., 2020) allows for variable amounts of work to be performed locally across devices, and relies on a proximal term in model align loss to help stabilize the method. Theoretically, it offers convergence guarantees under conditions of non-identical data distributions and variable device workloads.
@@
MOON (Li et al., 2021) is a model-contrastive FL framework that enhances local training by leveraging model representation similarities through contrastive learning at the model level.
@@
FedProto (Tan et al., 2022) is the first federated prototype learning framework for FL heterogeneity. Instead of exchanging gradients, clients and the server share abstract class prototypes. FedProto aggregates local prototypes and distributes global ones back to clients to regularize local model training, aiming to align local prototypes with global standards while minimizing local classification errors.
FedNH (Dai et al., 2023) addresses class imbalance by enhancing both personalization and generalization of local models. FedNH distributes class prototypes uniformly in the latent space, infusing class semantics to prevent prototype collapse and enhance model performance. This dual approach improves local models, boosting the generalization of the global model and thus refining personalized models.
FedTGP (Zhang et al., 2024a) unlikes conventional methods that aggregate prototypes via weighted averaging, FedTGP uses adaptive contrastive Learning to train global prototypes on the server, enhancing prototype separability and preserving semantic integrity.
As for (3), it is the core of OpenFGL as a comprehensive benchmark. To provide future FGL researchers with a comprehensive testing library and a user-friendly development framework, we conducted a thorough review of recent FGL studies, encompassing both the Graph-FL (GCFL+ (Xie et al., 2021), FedStar (Tan et al., 2023)) and Subgraph-FL (FedSage+ (Zhang et al., 2021), Fed-PUB (Baek et al., 2023), FedGTA (Li et al., 2023b), FGSSL (Huang et al., 2023), FedGL (Chen et al., 2024), AdaFGL (Li et al., 2024b), FGGP (Wan et al., 2024), FedTAD (Zhu et al., 2024), FedDEP (Zhang et al., 2024b)) scenarios, and comprehensively integrated them. The FGL baseline details implemented in our proposed OpenFGL are listed below:
@@
FedSage+ (Zhang et al., 2021) integrates node features, link structures, and labels using a GraphSage model and FedAvg across local subgraphs. FedSage+ extends this by training a generator to address missing links, enhancing model robustness and completeness.
@@
AdaFGL (Li et al., 2024b) introduces a two-step personalized approach that first aggregates multi-client models into a federated knowledge extractor during the final round at the server. Subsequently, each client undertakes personalized training utilizing the local subgraph and this federated extractor.
@@
FedDEP (Zhang et al., 2024b) leverages GNN embeddings for deep neighbor generation based on the FedSage+, employing efficient pseudo-FL for neighbor generation through embedding prototyping, and ensuring privacy protection via noiseless edge local DP. Meanwhile, it utilizes prototype representation technologies to further reduce communication costs.

### v2

@@ 4.4. Efficiency Evaluation
To answer Q8, we provide the efficiency reports in Fig. 2, Table 11, and Table 12, and obtain the following observations: (1) Prototype-based FedProto, FedTGP, and FGGP reduce communication costs but require extra computation, offsetting their runtime advantage. (2) Cross-client FedGL and FedSage+ suffer reduced efficiency due to delays from inter-client communication. (3) Decoupled AdaFGL maximizes local computation and minimizes communication costs, providing efficiency advantages. (4) In Graph-FL, models treat data samples independently, so even with large-scale graph samples, the limited size of each graph minimizes OOM or OOT issues. However, no single approach consistently achieves superior performance. (5) In our experiments, Param. refers to trainable parameters, Time to total runtime, and Com. to communication costs. FedSage+ and AdaFGL perform better but incur high Com. and computational complexity. Consequently, we conclude that C10: FGL algorithms leveraging prototypes and decoupled techniques (i.e., multi-client collaboration then local updates) demonstrate substantial potential in applications with stringent efficiency requirements (Heinbaugh et al., 2023; Li et al., 2024g).
4.5. FGL Guidance and OpenFGL Tutorial
In this section, we provide a detailed overview of OpenFGL: (1) Comprehensive evaluation of FGL methods to guide deployment (FGL Guidance); (2) User-friendly APIs to facilitate the reproduction and further development of FGL algorithms (OpenFGL Tutorial).
FGL Guidance. In Sec.4.1-Sec.4.4, we conduct a comprehensive empirical investigation of effectiveness, robustness, and efficiency, drawing 10 conclusions. These insights are crucial for selecting appropriate FGL algorithms in real-world applications. To provide a clearer presentation, we summarize the prevalent FGL algorithms in Table 13. Notably, scalability refers to the ability of a method to handle large-scale graphs without OOM or OOT. Regarding parallelism, we evaluate whether the method relies heavily on server-broadcasted information for local training. Minimal dependency allows for fewer communication and enables more independent parallel client training. Based on this, we observe that current FGL algorithms struggle to maintain consistent competitiveness across various requirements, highlighting that the field is still in its early stages with significant potential for future development.
OpenFGL Tutorial. We now present the algorithm design principles, which offer a unified API. It uses the FGLTrainer to manage client-server communication during each training round, aggregating messages via the message_pool variable. Users can customize the FGLClient and FGLServer to adjust message content and rules for specific algorithms. For clarity, we provide PyTorch-style implementations, exemplified by the FedAvg algorithm:
@@ 4.5. FGL Guidance and OpenFGL Tutorial
self.clients = load_client(args,...)
@@
(a) FGLTrainer. This class manages message and command flows between clients and a central server. In each training round, the trainer selects clients, updates the message_pool, and dispatches server messages. Clients process tasks locally and send updates to the server for global aggregation. Each round includes an evaluation phase. The implementation of FGLTrainer is shown in Algorithm 1.
@@
def __init__(self, args, client_id...):
@@
for (l_param, g_param) in zip(self.task.model.parameters(), self.message_pool[server][weight]):
@@
num_samples: self.task.num_samples,
weight: list(self.task.model.parameters())}
@@
(b) FedAvgClient. Users only need to customize: (1) Local execution, where FedAvg downloads the global model and performs local training; (2) Client-to-server messages, including the local sample count and model weights in FedAvg. The Pytorch-style implementation of FedAvgClient is provided in Algorithm 2.
@@
def __init__(self, args, message_pool...):
@@
# Receive messages (number of samples, local model weight) from each client stored in the message pool
@@
for it, client_id in enumerate(sampled_clients):
weight = self.message_pool[client_id][num_samples]/num_tot_samples (FedAvg_model_aggregation...)
@@
weight: list(self.task.model.parameters())}
Algorithm 3 OpenFGL-FGLServer: Pytorch style.
(c) FedAvgServer. Similar to FedAvgClient, users customize two modules: (1) Global execution (e.g., in FedAvg: weighted model averaging based on sample size); (2) Server-to-client messages (e.g., in FedAvg: global model weights). The Pytorch-style implementation of FedAvgServer is illustrated in Algorithm 3.
@@ 5. Conclusion and Future Directions
In this paper, we first present a comprehensive overview of the current research progress in the FGL field and the significant potential of this technology for deployment in graph-based database applications. Subsequently, we propose OpenFGL, a comprehensive FGL benchmark, which encompasses 18 recently proposed SOTA FGL algorithms and 42 datasets from 18 domains for 5 downstream tasks across 2 prevalent FGL scenarios. The goal of our work is to fairly examine the current state of FGL development and offer key insights for future research endeavors. Although FGL primarily serves downstream tasks in graph-based ML, it also holds significant potential for database applications. Specifically, it enables each client to generate high-quality graph embeddings (nodes, edges, and subgraphs) in a privacy-preserve and distributed manner. These embeddings provide rich semantic representations, which can be leveraged for efficient retrieval in vector databases. This perspective introduces new opportunities for applying FGL techniques to enhance the performance and scalability of graph-based databases.
Subsequently, we conduct extensive experiments aimed at unveiling the performance of FGL algorithms from 3 orthogonal perspectives: effectiveness, robustness, and efficiency. Our investigations reveal promising advancements achieved by FGL studies but also highlight their limitations, such as vulnerability in inadequate node descriptions, robustness, and scalability. To inspire future research, we combine experimental Conclusions to present the following significant challenges and promising directions.
@@
For Robustness, (3) Personalized FGL (C5 and C7). In real-world scenarios, the robustness of FGL against client-specific noise and low client participation in communication is essential. During federated training, these factors significantly impact the attainment of global consensus, thereby hindering high-quality supervision provided for local training. Fortunately, personalized techniques can leverage local reliable knowledge to establish unbiased global consensus and customize broadcasts for each client, guiding local updates. This mitigates optimization challenges arising from noise data and limited data sources. Consequently, personalized FGL algorithms emerge as an effective strategy for addressing these concerns. (4) Multi-client Collaboration FGL (C6 and C8). During our investigation, we found that promoting server-side multi-client collaboration can extract global insights from sparse data. Additionally, this collaborative approach can capture shared semantic knowledge across data domains to facilitate robust generalization. Therefore, there is an emergency demand for more efficient multi-client collaboration FGL algorithms. (5) Privacy-preserve FGL (C9). The goal of FGL is to achieve multi-client collaborative training in a privacy-preserving manner without direct data sharing. However, current FGL algorithms, in pursuit of superior performance, increasingly share local information, raising potential concerns. Therefore, the development of FGL algorithms with strict privacy requirements is imperative. Meanwhile, more privacy-preserving technologies should be considered for further explorations.
@@
FGL should establish federated collaboration standards for various graph types (e.g., directed, signed, hypergraphs, heterogeneous) and learning paradigms (e.g., unsupervised, few-shot, continual, unlearning) based on the data systems (Tian, 2023; Zhou et al., 2023; Besta et al., 2022). However, FGL remains a burgeoning field with numerous research gaps. Nevertheless, we are committed to continually enhancing OpenFGL to support future research endeavors. For instance, we are progressively refining the execution standards for federated heterogeneous graph learning. Notably, considering the space constraints and the need for a clear and reader-friendly presentation, we provide an overview and corresponding experimental results in  (Ope, 2024) (A.9).
@@ A.1. Dataset Description
In Graph-FL, we conduct experiments on the compounds networks (MUTAG, BZR, COX2, DHFR, PTC-MR, AIDS, NCI1, hERG, ogbg-molhiv, ogbg-molpcba) (Debnath et al., 1991; Sutherland et al., 2003; Helma et al., 2001; Riesen and Bunke, 2008; Wale et al., 2008; Gaulton et al., 2017; Hu et al., 2020b), protein networks (ENZYMES, DD, PROTEINS, ogbg-ppa) (Borgwardt et al., 2005; Dobson and Doig, 2003; Helma et al., 2001; Hu et al., 2020b), collaboration network (COLLAB) (Leskovec et al., 2005), movie networks (IMDB-BINARY, IMDB-MULTI) (Yanardag and Vishwanathan, 2015), super-pixel networks (MNISTSuperPixels) (Monti et al., 2017), point cloud networks (ShapeNet) (Yi et al., 2016), and yntax trees (ogbg-code2) (Hu et al., 2020b). In Graph-FL, node features generally include both node attributes and labels, which provide essential information for characterizing individual nodes within a graph. In most graph-level GNNs, these components are concatenated and subsequently fed into the model. As a result, we report the feature dimension as the combined size of both node attributes and labels. Furthermore, super-pixel networks provide an advanced approach to representing images as graphs, where nodes correspond to super-pixels, and edges capture spatial or semantic relationships between these regions. This graph-based representation effectively harnesses both the structural and semantic properties of the image, resulting in enhanced classification performance by moving beyond simple pixel-level analysis. A detailed description of the datasets used in our experiments is provided below:
@@
BZR (Sutherland et al., 2003) is a bioinformatics dataset aimed at predicting compound activity, with a primary focus on a diverse collection of benzimidazole compounds. This dataset is designed to provide information on the concentration required for each compound to inhibit the activity of specific biomolecules, thereby offering valuable insights into their effectiveness in modulating biological processes. Such data are critical for understanding the biochemical interactions of benzimidazole compounds and can serve as a resource for developing novel therapeutics by identifying promising candidates for further exploration in drug discovery.
COX2 (Sutherland et al., 2003) is a dataset focused on Cyclooxygenase-2, an enzyme that plays a crucial role in mediating inflammation and pain. The dataset is used to classify a variety of chemical compounds and predict their potential inhibitory potency against the COX-2 enzyme. Given that COX-2 is a key target in the development of anti-inflammatory drugs, this dataset is highly valuable for computational chemistry and drug discovery research, facilitating the identification of promising candidate compounds for therapeutic intervention in inflammatory conditions.
@@
ogbg-molhiv, ogbg-molpcba (Hu et al., 2020b) are molecular property prediction datasets of different sizes: ogbg-molhiv (small) and ogbg-molpcba (medium). Both are derived from MoleculeNet and are among its largest datasets. Molecules are pre-processed with RDKit, where each graph represents a molecule with nodes as atoms and edges as chemical bonds. Input node features are 9-dimensional, including atomic number, chirality, and other properties. Full feature descriptions are in the code. A script to convert SMILES strings into graph objects, which requires RDKit, can pre-process external molecular datasets for compatibility with OGB datasets.
ENZYMES (Borgwardt et al., 2005) is a comprehensive dataset consisting of 600 protein tertiary structures, meticulously curated from the BRENDA enzyme database. This dataset provides researchers with the opportunity to study the intricate and diverse structures of six distinct enzyme classes, enabling in-depth computational analysis and facilitating the development of advanced machine learning models for biological research. The rich structural diversity captured in ENZYMES makes it a valuable resource for tasks such as protein classification and function prediction, and other bioinformatics applications, ultimately contributing to a deeper understanding of enzyme mechanisms.
@@
ogbg-ppa (Hu et al., 2020b) comprises undirected protein association neighborhoods extracted from networks of 1,581 species, covering 37 taxonomic groups (e.g., mammals, bacteria, archaeans) across the tree of life. To construct these neighborhoods, 100 proteins were randomly selected from each species, and 2-hop neighborhoods were built. The center node was removed, and neighborhoods were sub-sampled to ensure fewer than 300 nodes. Nodes represent proteins, and edges denote biologically meaningful associations, with 7-dimensional features indicating the confidence (0-1) of specific associations, such as gene co-occurrence, gene fusion, and co-expression.
@@
ogbg-code2 (Hu et al., 2020b) comprises abstract syntax trees derived from approximately 450,000 Python method definitions. These methods are extracted from 13,587 repositories across popular GitHub projects. The dataset originates from GitHub CodeSearchNet, a collection of datasets and benchmarks for machine learning-based code retrieval. In ogbg-code2, we add an additional feature extraction step, including AST edges, AST nodes, and tokenized method names. This enables ogbg-code2 to capture the source code’s graph structure, beyond its token sequence representation.
@@ A.2. Baseline Description
Regarding (2), while some prevalent FL approaches (FedAvg (McMahan et al., 2017), FedProx (Li et al., 2020b), Scaffold (Karimireddy et al., 2020), MOON (Li et al., 2021a), and FedDC (Gao et al., 2022)) have demonstrated effectiveness in computer vision-based (CV-based) federated scenarios, recent FGL studies contend that they only achieve sub-optimal performance. This weakness is attributed to their failure to incorporate topological information during the collaborative optimization process of FGL. Therefore, we only implement prevalent FL as baselines to assist future FGL developers in evaluating the effectiveness of their proposed methods.
@@
Building upon this foundation, to provide effective baselines, we have particularly implemented prototype-based methods (FedProto (Tan et al., 2022), FedNH (Dai et al., 2023), FedTGP (Zhang et al., 2024b)). This has sparked a research trend in recent CV-based FL, as this technology can comprehensively address the aforementioned multi-level heterogeneity challenges. The prevalent graph-independent FL baseline details implemented in our proposed OpenFGL are listed below:
FedAvg (McMahan et al., 2017) serves as a foundational method in FL, enabling decentralized model training across diverse devices while preserving data privacy. Initiated by a central server that distributes a global model, clients independently execute local updates.
FedProx (Li et al., 2020b) allows for variable amounts of work to be performed locally across devices, and relies on a proximal term in model align loss to help stabilize the method. Theoretically, it offers convergence guarantees under conditions of non-identical data distributions and variable device workloads.
@@
MOON (Li et al., 2021a) is a model-contrastive FL framework that enhances local training by leveraging model representation similarities through contrastive learning at the model level.
@@
FedProto (Tan et al., 2022) is the first federated prototype learning framework for FL heterogeneity. Instead of exchanging gradients, clients and the server share abstract class prototypes. FedProto aggregates local prototypes and distributes global ones back to clients to regularize local model training.
FedNH (Dai et al., 2023) addresses class imbalance by enhancing both personalization and generalization of local models. FedNH distributes class prototypes uniformly in the latent space, infusing class semantics to prevent prototype collapse and enhance model performance. This dual approach improves local models.
FedTGP (Zhang et al., 2024b) unlikes conventional methods that aggregate prototypes via weighted averaging, FedTGP uses adaptive contrastive Learning to train global prototypes on the server, enhancing prototype separability and preserving semantic integrity.
As for (3), it is the core of OpenFGL as a comprehensive benchmark. To provide future FGL researchers with a comprehensive testing library and a user-friendly development framework, we conducted a thorough review of recent FGL studies, encompassing both the Graph-FL (GCFL+ (Xie et al., 2021), FedStar (Tan et al., 2023)) and Subgraph-FL (FedSage+ (Zhang et al., 2021), Fed-PUB (Baek et al., 2023), FedGTA (Li et al., 2023b), FGSSL (Huang et al., 2023), FedGL (Chen et al., 2024), AdaFGL (Li et al., 2024e), FGGP (Wan et al., 2024), FedTAD (Zhu et al., 2024), FedDEP (Zhang et al., 2024c)) scenarios, and comprehensively integrated them. The FGL baseline details implemented in our proposed OpenFGL are listed below:
@@
FedSage+ (Zhang et al., 2021) integrates node features, link structures, and labels by employing a GraphSage model in conjunction with FedAvg to facilitate federated learning over local subgraphs. To further improve performance, FedSage+ introduces a generator to address missing links within the graph structure, thereby enhancing the model’s robustness against incomplete data and ensuring a more comprehensive representation of graph relationships. This approach ultimately strengthens the ability of the model to generalize effectively across different clients in a federated learning setting.
@@
AdaFGL (Li et al., 2024e) introduces a two-step personalized approach. In the first step, multi-client models are aggregated into a federated knowledge extractor at the server during the final communication round. In the second step, each client conducts personalized training using its local subgraph along with the federated knowledge extractor, allowing for more tailored model adaptation to local data characteristics while benefiting from collective knowledge.
@@
FedDEP (Zhang et al., 2024c) leverages GNN embeddings for deep neighbor generation based on the FedSage+, employing efficient pseudo-FL for neighbor generation through embedding prototyping, and ensuring privacy protection via noiseless edge local DP. Meanwhile, it utilizes prototype representation technologies to further reduce communication costs.